# Estimating Machine Translation Difficulty

### v1

@@ 6 Creating Difficult Benchmarks
In this section, we apply the top-performing difficulty estimators for creating difficult machine translation benchmarks. First, we show that the test set employed at the WMT 24 general MT shared task Kocmi et al. (2023) might be too easy for current MT models. Then, we define the task of selecting a difficult subset of samples from a given dataset and we evaluate our estimators.
@@
In Figure˜1, we show the distribution of scores assigned by human annotators to the translations of the WMT 2024 general MT shared task (Kocmi et al., 2024a). Notably, the best model for each language attains almost always 90 to 100 ESA points, and so does at least one system for each input text. This is particularly concerning for the English-to-Spanish translation direction, where top systems barely made any errors. These findings highlight the need to create more difficult benchmarks beyond randomly sampling texts. In contrast, the rest of this section uses difficulty estimators to samples difficult texts out of a larger collection.
@@ 6.1 Setup
Given a large set of source texts $\mathcal{X}$, we aim to extract a subset $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ of maximum difficulty of size $|\mathcal{X}^{\prime}|=B$. In practice, we select $B$ texts with the highest difficulty as determined by the top-performing difficulty estimators. As per Section˜5.2, we exclude internal artificial crowd-based estimators, as they would unfairly bias the benchmark against the evaluated MT models.
As in previous experiments, we use human translation quality annotations to assess the effectiveness of our estimators. We again rely on the WMT 24 test set, which we use as $\mathcal{X}$ (see Section˜5 for further details). Specifically, we focus on its English source texts, which were translated into Czech, Spanish, Hindi, Icelandic, Japanese, Russian, Ukrainian, and Chinese, as well as its Czech sources translated into Ukrainian.
We emphasize that we select $B$ source texts from $\mathcal{X}$. Thus, regardless of the target languages, any $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ includes only English and Czech source texts, while target languages are only considered for evaluation purposes.
@@
We assign a single difficulty score $\hat{d}(x)$ to each sample $x\in\mathcal{X}$. This is straightforward when using text-only estimators such as heuristics and learned methods, which do not use additional information aside from the given text. For external artificial crowd-based methods, we assign to each text $x$ the average estimated quality score of its translations, averaging across both the MT models and target languages. Then, we extract $\mathcal{X}^{\prime}$ as follows:
$\hat{\mathcal{X}}^{\prime}=\underset{x\in\mathcal{X}}{\mathrm{arg\,top}\text{-\hskip-1.00006pt}B}\,\,\,\hat{d}(x)$ (2)
@@
One goal of constructing a difficult benchmark is to identify samples where contemporary models still struggle, in order to expose their shortcomings and guide improvements in future iterations. Therefore, we evaluate the usefulness of difficulty estimators based on the drop in the average human score obtained by the models’ translations on the test set. As additional information, we also report the proportion of “perfect” outputs (those that received a full score of 100/100 ESA points from human annotators) that remain in the subsampled test set. The exact formulas for these measures are provided in Appendix˜C.
@@ 6.2 Results
We extract several $\mathcal{X}^{\prime}\in\mathcal{X}$ by varying the size of the subsample, and report the curves of the Average Score and %Perfect measures in Figure˜2. Unsurprisingly, the Oracle creates a bound on the best achievable performance, though on %Perfect it is matched closely by the dedicated predictor Sentinel. Both heuristic-based and LLM-as-a-Judge-based methods are very close in performance to random subset selection. The Oracle is worse when measured by %Perfect because it is optimized towards the lowest average score and not proportion of lowest %Perfect. More detailed quantitative results for the 25% test set subsample scenario, including per-domain breakdowns, are reported in Appendix˜F. Returning to the original lament of existing testsets being easy (Figure˜1), in Appendix Figure˜5 we show how the score distribution changes upon selecting difficult texts.
@@
Table 3: Pearson correlations between selectors and variables of interest (source length, source number of errors per word, output diversity measured by pairwise embeddings inner product and chrF, and proportion of unique outputs). For all methods, lower values indicate more difficult source texts, so negative correlation implies stronger positive connection between difficulty and the target variable. See Appendix Figure˜6 for detailed visualization.
@@ 6.3 Potential Pitfalls of Selecting by Difficulty
In practice, selecting samples by anything other than random sampling harbors unexpected dangers. For example, the texts selected by a difficulty estimator might be grammatically incorrect. We now investigate potential pitfalls one might encounter when using difficulty as the subsampling criterion. Specifically, we focus on:
  • Source length: Difficulty estimators might over-rely on text length. While longer texts are more difficult to translate compared to shorter ones, their translations are also more expensive to human-evaluate.
  • Source errors: Translating incomprehensible source texts is naturally difficult. Nevertheless, these might be undesirable to be in the test set.
  • Output diversity: Source texts that lead to diverse machine translation outputs are more desirable, as they help distinguish between models. See Appendix˜C for implementation details.
We present the correlation between difficulty estimators’ predictions and these variables of interest in Table˜3 and Appendix Figure˜6. Given that we defined difficulty using translation quality, lower estimators’ scores indicate higher difficulty.
As expected, all estimators show a strong negative correlation with source length, indicating they are all biased toward selecting longer outputs. On the contrary, this does not seem to be the case for Source errors, meaning that texts containing many errors seem not to be prioritized by difficulty estimators. Finally, our results suggest that all difficulty estimators select source texts that lead to more diverse and unique outputs.
The original dataset $\mathcal{X}$ comes from different domains and subdomains. A legitimate concern is whether selecting by difficulty reduces the diversity of the subset, which we measure with embeddings and vector space distribution. Appendix Figure˜4 shows that with larger subsets, the space becomes more densely populated. Naturally, random selection is a good baseline for creating a diverse subset. On the other hand, all other difficulty estimation methods, including the Oracle, lead to even denser space, implying lower diversity. If test set diversity is desired, explicit measures need to be adopted, such as first identifying diverse clusters and selecting by difficulty estimation within each cluster.
Evaluating difficulty estimation methods on the WMT 2024 has a limitation. This test set was already downsampled from a larger batch and vetted by humans, making the distribution of phenomena artificial. To mitigate this issue, we conduct this analysis using a larger batch of data and report results in Appendix Figure˜4.
@@
Texts from the WMT24 dataset of English sources with Sentinel-src decisions on difficulty (easy=lowest difficulty, difficulty=highest difficulty in each length bin).
@@ 6.4 Qualitative Analysis
The previous high-level analyses revealed the Sentinel-src is an apt selector, with less bias risk than internal artificial crowd and outperforming other methods. To further verify this model, we manually inspected 200 source texts, half deemed easy and half difficult by Sentinel-src, separated each in 10 buckets by length. In general, the notion of difficulty by Sentinel-src aligns well with human judgments of translation difficulty. For instance, difficult segments tend to use idiomatic expressions (Table˜4.1). Segments that are not full sentences, such as headlines also receive high difficulty (Table˜4.2), which is aligned with human perception. The same holds true for multi-sentence segments with indirect speech (Table˜4.3). On the other hand, segments with single words, simple sentence structure, or simply multiple joined simple sentences, are perceived as easy (Table˜4.4 to Table˜4.6).
@@ 7 Conclusion
In this work, we formally define the task of translation difficulty estimation and introduce a dedicated evaluation measure to assess the performance of difficulty estimators. We then conduct a comprehensive evaluation of both existing and newly proposed estimators, finding that traditional heuristic-based methods are outperformed by alternative approaches. Notably, we find that LLM-as-a-judge approaches perform poorly, being surpassed by all methods explicitly trained for the task. Among these, we identify Sentinel-src-24 as the current state-of-the-art for this task and recommend it for downstream applications. To this end, we examine its performance alongside other estimators in the difficult benchmark creation task, demonstrating these models effectively identify difficult samples where current MT models underperform. Sentinel-src-24 remains the top-performing method in this downstream application as well. Inspired by this, we develop Sentinel-src-25 by incorporating additional data into Sentinel-src-24’s training pipeline, and release both models publicly. Finally, we conduct a qualitative analysis of Sentinel-src-25’s predictions, offering intuitive insights into the types of texts it deems difficult.
@@ Appendix A Kendall $\bm{\tau}_{\bm{b}}$
$\tau_{b}=\frac{C-D}{\sqrt{(C+D+T_{h})(C+D+T_{m})}}$ (3)
Where $C$ and $D$ are the number of concordant and discordant pairs, respectively, between $R$ and $\hat{R}$, and $T_{h}$ and $T_{g}$ are the number of tied pairs in the hypothesis ranking and the ground truth ranking, respectively. Specifically, the pair $(i,j)$ is concordant if $\mathrm{sign}(r_{i}-r_{j})=\mathrm{sign}(\hat{r}_{i},\hat{r}_{j})$ and discordant if $\mathrm{sign}(r_{i}-r_{j})\neq\mathrm{sign}(\hat{r}_{i},\hat{r}_{j})$. Finally, the pair $(i,j)$ is tied in the hypothesis ranking or the ground truth ranking if $\hat{r}_{i}=\hat{r}_{j}$ or $r_{i}=r_{j}$, respectively.
We chose Kendall’s $\tau_{b}$ over alternatives like Pearson correlation such that it is not affected by outliers. The $\tau_{b}$ is also a standard in machine translation meta-evaluation (Deutsch et al., 2023) as it handles ties better than Spearman correlation.
Appendix B Training Sentinel-MQM-24 and Sentinel-MQM-25
@@
Our new learned difficulty estimation models, Sentinel-MQM-24 and Sentinel-MQM-25, follow the same architecture and training pipeline used for the Sentinel-MQM model introduced by Perrella et al. (2024). Both models are based on XLM-RoBERTa large as the backbone encoder, followed by a multi-layer feedforward network on top of the [CLS] token. They are trained to minimize the Mean Squared Error (MSE) between predicted and human scalar scores.
@@
We adopt the same two-stage training approach as the Sentinel-MQM model. In the first stage, the model is trained on Direct Assessment (DA, Graham et al., 2013) data. In the second stage, it is fine-tuned on MQM annotations. The key differences between our models and Sentinel-MQM lie in the training data used at each stage.
@@
  • Stage 1: DA training. For Sentinel-MQM-24, we extend the DA training data used by Perrella et al. (2024) by including annotations from WMT 21 Wenzek et al. (2021), as well as DA+SQM annotations from WMT 22 Kocmi et al. (2022) and WMT 23 Kocmi et al. (2023). Sentinel-MQM-25 further includes the ESA annotations from WMT 24.
@@
  • Stage 2: MQM fine-tuning. In this phase, we expand the MQM training set by adding MQM annotations from WMT 23 Freitag et al. (2023). Unlike the Sentinel-MQM training pipeline, we do not average multiple scores per translation. Instead, we include all available annotations as individual training instances, preserving variability across raters. This applies to WMT 20 and WMT 22 MQM datasets, which include three human scores per translation Freitag et al. (2021); Riley et al. (2024). Similarly to the first training stage on DA, in the case of the Sentinel-MQM-25 model, we also include MQM annotations from WMT 24.
@@
Following the approach of Perrella et al. (2024), we treat each pair consisting of a source text segment and its associated human score as an independent training instance. Since human scores are assigned to individual translations, multiple annotations may exist for the same source text. We do not combine these scores in any way but include them all in the training data for both DA and MQM stages. Training hyperparameters match those used by Perrella et al. (2024) for Sentinel-MQM. All models were trained using a single NVIDIA GeForce RTX 4090 GPU. The estimated training time is approximately three GPU hours for the first (DA) stage and one GPU hour for the second (MQM) fine-tuning stage. These estimates apply to both Sentinel-MQM-24 and Sentinel-MQM-25.
@@ Appendix C Implementation Details
  • CommandA:  [^11] B-parameter LLM for real-world enterprise use cases Cohere et al. (2025).
@@
Figure 3: Average similarity between two closest (left) and any two (right) source texts in $\mathcal{X}$ based on embeddings and cosine similarity. The curves go up because the vector space saturates and nearest neighbours become closer. Random selection shows 99% confidence t-test interval from 10 runs.
Figure 4: In contrast to Figure˜4, we have collected contemporary news articles (40k segments through crawling, Ba
textasciitilde nón et al., 2020
) to evaluate how our difficulty sampling would perform in a real world. Average similarity between two closest (left) and any two (right) source texts based on embeddings and cosine similarity in $\mathcal{X}$ on raw 40k English segments (not WMT24). The curves go up because the vector space saturates and nearest neighbours become closer. Random selection shows 99% confidence t-test interval from 10 runs. LLM-as-a-Judge and Artificial crowd were not included due to compute costs. Oracle is not present due to the absence of model outputs and human scores.
@@
Sentinel-MQM-24 [Uncaptioned image] [Uncaptioned image] $0.246$
Sentinel-MQM [Uncaptioned image] [Uncaptioned image] $0.235$
@@
Artificial Crowd
Internal (XCOMET-QE-XXL) [Uncaptioned image] [Uncaptioned image] $\mathbf{0.278}$
Internal (MetricX-24-Hybrid-QE-XXL) [Uncaptioned image] [Uncaptioned image] $0.248$
External (XCOMET-QE-XXL) [Uncaptioned image] [Uncaptioned image] $0.207$
External (MetricX-24-Hybrid-QE-XXL) [Uncaptioned image] [Uncaptioned image] $0.185$

### v2

@@ 6 Creating Difficult Benchmarks
In this section, we use the top-performing difficulty estimators to create difficult machine translation benchmarks. First, we show that the test set employed at the WMT 2024 General MT shared task Kocmi et al. (2023) is too easy for current MT models. Then, we define the task of selecting a difficult subset of samples from a given dataset, and evaluate our estimators.
@@
In Figure˜1, we report the distribution of scores assigned by human annotators to the translations of the WMT 2024 General MT shared task (Kocmi et al., 2024a). Notably, the best model for each language attains almost always 90 to 100 ESA points, and so does at least one system for each input text. This is particularly concerning for the English-to-Spanish translation direction, where top systems barely made any errors. These findings highlight the need to create more difficult benchmarks. To do this, we use difficulty estimators to sample difficult texts out of a larger collection.
@@ 6.1 Setup
Given a large set of source texts $\mathcal{X}$, we aim to extract a subset $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ of maximum difficulty of size $|\mathcal{X}^{\prime}|=B$. In practice, we select $B$ texts with the highest difficulty as determined by the top-performing difficulty estimators. As discussed in Section˜5.2, we exclude True Crowd estimators.
We again rely on the WMT 24 test set, which we use as $\mathcal{X}$ (see Section˜5 for further details). Specifically, we focus on its English source texts, which were translated into Czech, Spanish, Hindi, Icelandic, Japanese, Russian, Ukrainian, and Chinese, as well as its Czech sources translated into Ukrainian. Accordingly, since we select $B$ source texts from $\mathcal{X}$, any subset $\mathcal{X}^{\prime}\subseteq\mathcal{X}$ necessarily contains only English and Czech source texts, while target languages are considered solely for evaluation purposes.
@@
We assign a single difficulty score $\hat{d}(x)$ to each sample $x\in\mathcal{X}$. For source text-only difficulty estimators, such as heuristics and learned methods, this is straightforward, as they rely solely on the given source text. Instead, for Artificial Crowd methods, we assign to each text $x$ the average quality score of its translations, estimated using XCOMET, averaging across both the employed MT models and the target languages. Finally, we construct $\mathcal{X}^{\prime}$ by selecting the $B$ most difficult source texts.
@@
One goal of constructing a difficult benchmark is to identify samples where contemporary models still struggle, in order to expose their shortcomings and guide improvements in future iterations. Therefore, we evaluate the usefulness of difficulty estimators based on the drop in the average human score obtained by the models’ translations on the test set. As additional information, we also report the proportion of “perfect” outputs (i.e., those that received a full score of 100/100 ESA points from human annotators) that remain in the subsampled test set. The exact formulas for these measures are provided in Appendix˜C.
@@ 6.2 Results
We extract several $\mathcal{X}^{\prime}\in\mathcal{X}$ by varying the size of the subsample, and report the curves of the Average Score and %Perfect measures in Figure˜2. First, we wish to highlight that the oracles serve as a performance upper-bound only in terms of Average Score, and not in terms of %Perfect, because they are designed to select the sources with the lowest average score, rather than the lowest %Perfect. In this respect, oracle-src selects the sources with the lowest average difficulty score across models and target languages; instead, oracle-tgt selects a different source text for each target language, averaging difficulty scores only across MT models. As we can see from Figure˜2, text length heuristics and LLM-as-a-Judge-based methods show very close performance to random subset selection, especially in terms of Average Score. Instead, Sentinel-src-24 and Artificial Crowd perform closer to the oracles, even surpassing oracle-src in terms of %Perfect. In Appendix˜F, we report detailed quantitative results for the scenario where we subsample 25% of the test set, including per-domain performance breakdowns.
Additionally, returning to the original lament of existing test sets being too easy (Figure˜1), in Appendix Figure˜6 we show how the score distribution changes when selecting difficult texts.
@@
Table 3: Pearson correlations between difficulty estimators and variables of interest (source length, number of errors per source word, output diversity, and proportion of unique outputs). All estimators assign lower values to more difficult source texts. Therefore, negative correlation indicates a positive correlation between difficulty and the variable of interest. See Appendix Figure˜7 for detailed visualization and Appendix˜C for implementation details.
@@ 6.3 Potential Pitfalls of Selecting by Difficulty
Selecting samples by anything other than random sampling may harbor unexpected dangers. For example, the texts selected by a difficulty estimator might be grammatically incorrect or poorly formed. Here, we investigate potential pitfalls one might encounter when using difficulty as a subsampling criterion. Specifically, we focus on:
  • Source length: Longer texts are more difficult to translate compared to shorter ones. We are interested in quantifying the extent to which difficulty estimators rely on text length.
  • Source errors: Translating incomprehensible source texts is naturally difficult. Nevertheless, it might be undesirable to create test sets containing many garbled sources.
  • Output diversity: When creating a benchmark, source texts that lead to more diverse outputs are more desirable, as they help distinguish between models. See Appendix˜C for implementation details.
We present the correlation between difficulty estimators’ predictions and these variables of interest in Table˜3 and Appendix Figure˜7. We wish to remind the reader that we defined difficulty using translation quality, meaning that lower estimators’ scores indicate higher difficulty. Therefore, a negative correlation with difficulty scores should be interpreted as a strong correlation with the concept of difficulty, as estimated by our models.
As expected, all estimators show a strong negative correlation with source length, indicating they are all biased toward selecting longer outputs. On the contrary, this does not seem to be the case for Source errors, suggesting that our difficulty estimators do not prioritize texts containing many errors. Finally, our results suggest that Sentinel-src-24 and Artificial Crowd select source texts that lead to more diverse outputs.
[…]
Bias toward MT models.
[…]
Overall, the mean Kendall $\tau_{b}$ correlations fall within a narrow range and have comparable standard deviations. This suggests that the notion of difficulty captured by both estimators aligns relatively uniformly with MT models’ perceived difficulty. However, we note two exceptions: ONLINE-B exhibits a higher correlation than average with Sentinel-src-24, whereas Gemini-1.5-Pro shows a lower average correlation with both estimators. While this is a preliminary investigation, this finding suggests that a benchmark created using these estimators could be disproportionately difficult for ONLINE-B and easier for Gemini-1.5-Pro. We leave for future work a deeper investigation into whether this discrepancy stems from a bias in the difficulty estimations or simply from variance, especially given that the correlations were averaged over a limited number of translation directions.
@@
The most difficult and easiest English source texts from the WMT24 dataset, as selected by Sentinel-src.
@@ 6.4 Qualitative Analysis
We manually inspected 200 source texts, half of which were deemed easy and half difficult by Sentinel-src-25, and we separated them into 10 length-based buckets. In general, difficulty levels assigned by Sentinel-src-25 align well with human perception of difficulty. Indeed, we find that difficult segments often contain complex constructions (Table˜4.1), consist of incomplete sentences, such as headlines (Table˜4.2), or include indirect speech (Table˜4.3). On the other hand, the segments classified as easy by Sentinel-src-25 are typically single words, have simple sentence structures, or are concatenations of short, simple sentences (Table˜4.4 to Table˜4.6).
@@ 7 Conclusion
In this work, we formally define the task of translation difficulty estimation and introduce the Difficulty Estimation Correlation (DEC), a dedicated measure to evaluate the performance of difficulty estimators. We conduct a comprehensive evaluation of both existing and newly proposed estimators, finding that models explicitly trained for the task significantly outperform traditional, heuristic-based methods and LLM-as-a-judge approaches.
Our analysis identifies Sentinel-src-24 as the current state-of-the-art in translation difficulty estimation. We further validate the performance of difficulty estimators in the downstream task of creating difficult benchmarks, demonstrating that they successfully identify samples where modern MT models underperform. Also in this downstream task, Sentinel-src-24 remains the top-performing method. Building on these findings, we develop Sentinel-src-25 by incorporating additional data into the training pipeline of Sentinel-src-24, and release both models publicly. Finally, we conduct a qualitative analysis of Sentinel-src-25’s predictions, offering intuitive insights into the types of texts it deems difficult.
@@
Using the WMT 2024 test set to analyze difficulty estimators.
In Section˜6.3, we investigated potential concerns of subsampling large data sets using our difficulty estimators. However, to do this, we used the WMT 2024 test set, which has a limitation. The sources contained in this test set were vetted by humans, making the distribution of the phenomena we investigate artificial. To mitigate this issue, we conduct the same analysis using a larger batch of data and report results in Appendix Figure˜5.
@@ Appendix A Kendall $\bm{\tau}_{\bm{b}}$
$\tau_{b}\;=\;\frac{C-D}{\sqrt{(C+D+T_{g})(C+D+T_{h})}}\,.$ (2)
Here $C$ and $D$ are the numbers of concordant and discordant pairs when comparing gold scores $r_{i}$ with hypothesis scores $\hat{r}_{i}$: a pair $(i,j)$ is concordant if $(r_{i}-r_{j})(\hat{r}_{i}-\hat{r}_{j})>0$ and discordant if $(r_{i}-r_{j})(\hat{r}_{i}-\hat{r}_{j})<0$. $T_{h}$ counts pairs tied only in the hypothesis ($\hat{r}_{i}=\hat{r}_{j}$ and $r_{i}\neq r_{j}$), and $T_{g}$ counts pairs tied only in the gold ($r_{i}=r_{j}$ and $\hat{r}_{i}\neq\hat{r}_{j}$); pairs tied in both rankings are ignored.
We use Kendall’s $\tau_{b}$ rather than Pearson correlation because $\tau_{b}$ evaluates relative order (ranks) and is therefore more robust to scale differences and outliers. This is important in our setting: as visible in Figure˜1, most WMT24 segment-level scores lie in a narrow band (e.g., $[90,100]$), which magnifies the effect of outliers on Pearson, whereas $\tau_{b}$ depends only on pairwise rankings. This behavior is also discussed by Mathur et al. (2020).
Appendix B Training Sentinel-src-24 and Sentinel-src-25
@@
Our new learned difficulty estimation models, Sentinel-src-24 and Sentinel-src-25, follow the same architecture and training pipeline used for the Sentinel-src model introduced by Perrella et al. (2024). Both models are based on XLM-RoBERTa large as the backbone encoder, followed by a multi-layer feedforward network on top of the [CLS] token. They are trained to minimize the Mean Squared Error (MSE) between predicted and human scalar scores.
@@
We adopt the same two-stage training approach as the Sentinel-src model. In the first stage, the model is trained on Direct Assessment (DA, Graham et al., 2013) data. In the second stage, it is fine-tuned on MQM annotations. The key differences between our models and Sentinel-src lie in the training data used at each stage.
@@
  • Stage 1: DA training. For Sentinel-src-24, we extend the DA training data used by Perrella et al. (2024) by including annotations from WMT 21 Wenzek et al. (2021), as well as DA+SQM annotations from WMT 22 Kocmi et al. (2022) and WMT 23 Kocmi et al. (2023). Sentinel-src-25 further includes the ESA annotations from WMT 24. For both model versions, we also incorporate MLQE-PE data Fomicheva et al. (2022) in the training set for this stage.
@@
  • Stage 2: MQM fine-tuning. In this phase, we expand the MQM training set by adding MQM annotations from WMT 23 Freitag et al. (2023). Unlike the Sentinel-src training pipeline, we do not average multiple scores per translation. Instead, we include all available annotations as individual training instances, preserving variability across raters. This applies to WMT 20 and WMT 22 MQM datasets, which include three human scores per translation Freitag et al. (2021); Riley et al. (2024). Similarly to the first training stage on DA, in the case of the Sentinel-src-25 model, we also include MQM annotations from WMT 24.
@@
Following the approach of Perrella et al. (2024), we treat each pair consisting of a source text segment and its associated human score as an independent training instance. Since human scores are assigned to individual translations, multiple annotations may exist for the same source text. We do not combine these scores in any way but include them all in the training data for both DA and MQM stages. Training hyperparameters match those used by Perrella et al. (2024) for Sentinel-src. All models were trained using a single NVIDIA GeForce RTX 4090 GPU. The estimated training time is approximately three GPU hours for the first (DA) stage and one GPU hour for the second (MQM) fine-tuning stage. These estimates apply to both Sentinel-src-24 and Sentinel-src-25.
@@ Appendix C Implementation Details
  • For the word rarity heuristic, we compute word frequencies using the wordfreq Python library Speer (2022).
  • For the syntactic complexity heuristic and text length, we obtain dependency trees and corresponding tokens using spaCy. Specifically, we use language-specific pipelines: i) en_core_web_sm for English, ii) ja_core_news_sm for Japanese, and iii) spacy_udpipe for Czech.
  • For the output diversity assessment (Section 6.3), we compute multilingual sentence embeddings from the source texts using sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 (Reimers and Gurevych, 2019). Specifically, for each pair of translations, we measure the inner product between their multilingual sentence embeddings and their chrF score computed one against the other.
@@
  • CommandA:  [^11] B-parameter LLM for real-world enterprise use cases Cohere Team (2025).
@@
Figure 4: Average similarity between two closest (left) and any two (right) source texts in $\mathcal{X}$ based on embeddings and cosine similarity. The left curves go up because the vector space saturates and nearest neighbours become closer. Random selection shows 99% confidence t-test interval from 10 runs.
Figure 5: In contrast to Figure˜5, we have collected contemporary news articles (40k segments through crawling, Bañón et al., 2020) to evaluate how our difficulty sampling would perform in a real world. Average similarity between two closest (left) and any two (right) source texts based on embeddings and cosine similarity in $\mathcal{X}$ on raw 40k English segments (not WMT24). The left curves go up because the vector space saturates and nearest neighbours become closer. Random selection shows 99% confidence t-test interval from 10 runs. LLM-as-a-Judge and Artificial crowd were not included due to compute costs. Oracle is not present due to the absence of model outputs and human scores.
@@
Sentinel-src-24 [Uncaptioned image] [Uncaptioned image] $0.246$
Sentinel-src [Uncaptioned image] [Uncaptioned image] $0.235$
@@
Crowd Based
True (XCOMET-QE-XXL) [Uncaptioned image] [Uncaptioned image] $\mathbf{0.278}$
True (MetricX-24-Hybrid-QE-XXL) [Uncaptioned image] [Uncaptioned image] $0.248$
Artificial (XCOMET-QE-XXL) [Uncaptioned image] [Uncaptioned image] $0.207$
Artificial (MetricX-24-Hybrid-QE-XXL) [Uncaptioned image] [Uncaptioned image] $0.185$